I need your help. I am using Excel for some problems.
Enter your coefficients and formula for x. Rearrange the equation in the form Ax + B = C.
22x + 60 = -11x - 6
-7 (x - 4) -11x - 6 = 2x
(3/2) = x + (2/5)
Can you explain??

Answer by stanbon(75887)

Rearrange the equation in the form Ax + B = C.
22x + 60 = -11x - 6
33x + 60 = -6
-----------------------
-7 (x - 4) -11x - 6 = 2x
-7x+28 -11x -6 = 2x
20x + 6 = 28
-------------------
(3/2) = x + (2/5)
x + +(2/5) = (3/2)
10x + 4 = 15
===============
Cheers,
Stan H.